OpenStudy (timaashorty):

Can anyone please help me with this question: Using Matrices to rotate the vector V= (-4,1) = by 270°. Write the resulting vector in component form.

OpenStudy (timaashorty):

@wio Can you pls help me?

OpenStudy (anonymous):

Do you remember the rotation matrix?

OpenStudy (timaashorty):

kind of

OpenStudy (anonymous):

\[ \begin{bmatrix} \cos\theta & -\sin\theta \\ \sin\theta & \cos\theta \end{bmatrix} \]Multiply this by a vector to rotate it \(\theta\) degrees.

OpenStudy (timaashorty):

can you show me how to do that?

OpenStudy (anonymous):

First, \(\theta = 270^\circ \) so: \[ \begin{bmatrix} \cos(270^\circ) & -\sin(270^\circ)\\ \sin(270^\circ) & \cos(270^\circ) \end{bmatrix} = \begin{bmatrix} 0 & 1\\ -1 & 0 \end{bmatrix} \]

OpenStudy (anonymous):

Multiply \[ \begin{bmatrix} 0 & 1\\ -1 & 0 \end{bmatrix} \begin{bmatrix} -4\\ 1 \end{bmatrix} \]

OpenStudy (timaashorty):

Im sorry I left and took long. My computer just crashed. anyways it'll be: [1] [4]

OpenStudy (timaashorty):

IS that it or do i need to do more steps?

OpenStudy (timaashorty):

@wio

OpenStudy (anonymous):

Yeah, that is the answer. Just graph to check.

OpenStudy (timaashorty):

Thank you so much wio!